OpenStudy (anonymous):

how long is a guy wire reaching from the top of a 6ft pole, to a point on the ground, 4ft from the pole?

OpenStudy (matt101):

Pythagorean Theorem: 6^2 + 4^2 = x^2 36 + 16 = x^2 52 = x^2 x ~ 7.21 ft